OpenStudy (anonymous):

who ever helps me gets a medal i need help What is the solution to r over three is equal to one over three? one over twelve twelve over one three over four four over three

OpenStudy (charlotte123):

So: r/3 = 1/3?

OpenStudy (anonymous):

well thats not one of the answers but i know that you are trying

OpenStudy (charlotte123):

I know xD Just wanted to make sure that that was It before I help you :)

OpenStudy (anonymous):

oh ok then that is correct

OpenStudy (charlotte123):

Hmm........ r would have to be 1 o.o Unless If that 1/3 was suppose to be: 1 1/3

OpenStudy (charlotte123):

Does that seem to be the case?

OpenStudy (anonymous):

you sound like bame every time you talk or you talk like you are just awsome so i love you lol

OpenStudy (charlotte123):

*supposedly My bad hueheuhe

OpenStudy (charlotte123):

Oh me oh my o.o Why thank you xD

OpenStudy (anonymous):

you are just outstanding

OpenStudy (charlotte123):

Now now now euheuheuehuehhe ^///////^

OpenStudy (charlotte123):

Remember to concentrate you fool @charlotte123

OpenStudy (charlotte123):

=.= @charlotte123

OpenStudy (charlotte123):

My bad my bad So dear @Blacktears21 what seems to be the case?Was the problem r/3 = 1/3 or r/3 = 1 1/3?:)

OpenStudy (anonymous):

i guess that does not seem to be the case?

OpenStudy (anonymous):

never mind i think i found the answer but i have a different question

OpenStudy (anonymous):

lets go

OpenStudy (charlotte123):

Hmm........ We seem to have a problem D: You see: 1/3 = 0.33333333333..... 1/12/3 = 0.02777777777..... 12/1/3 = 4 3/4/3 = 0.25 4/3/3 = 0.44444444....... None of them match what 1/3 Is = to :(

OpenStudy (charlotte123):

Okey dokey!

OpenStudy (charlotte123):

Go on @Blacktears21 :) Post away!

OpenStudy (charlotte123):

Ah my bad xD I see It now LOL